### v1

# When is the canonical conductor minimal?

Abstract.
For a one dimensional analytically unramified local domain $R$, the blowup algebra of the canonical ideal is a birational extension. The conductor of this extension always contains the conductor of $R$. We study the case when there is equality and we classify numerical semigroups with this property.
Key words and phrases:
almost Gorenstein rings, nearly Gorenstein rings, far-flung Gorenstein rings, trace ideals, conductor ideals
2020 Mathematics Subject Classification:
13C05, 13C14, 13G05, 13H10
1. Introduction
Gorenstein rings are ubiquitous as Hyman Bass pointed out over half a century ago [Bas63]. They possess beautiful symmetry properties and they are a special family inside the broader family of Cohen-Macaulay rings. In the last few decades, there has been a growing interest in understanding Cohen-Macaulay local rings which are close to being Gorenstein. This has started with the introduction of one dimensional almost Gorenstein rings by Barucci-Fröberg [BF97] and continued with the work of Goto-Matsuoka-Phuong [GMP13] and Goto-Takahashi-Taniguchi [GTT15] in higher dimensions. Later in [HHS19], Herzog-Hibi-Stamate introduced nearly Gorenstein rings and this is still a very active research are in commutative algebra. On the other extreme, Herzog-Kumashiro-Stamate [HKS23] considered far-flung-Gorenstein rings which are one dimensional rings that are not close to being Gorenstein.
The main actors in this type of study are trace ideals. Nearly Gorenstein rings and far-flung Gorenstein rings are defined by imposing conditions on the trace of the canonical module. Dao-Maitra-Sridhar showed in [DMS23, Proposition 7.1] that in dimension one, the almost Gorenstein property can also be seen by a condition on the trace of the blowup algebra of the canonical module. More precisely, let $R$ be a one dimensional generically Gorenstein local ring with canonical ideal $\omega$ and maximal ideal $\mathfrak{m}$. Denote by $\operatorname{\mathrm{tr}}(\omega)$ the trace of the canonical module (see Definition 2.1) and by $b(\omega)$ the conductor of the blowup algebra $B(\omega)$ of the canonical module (see Remark 2.2 and Remark 2.7). Then, under mild conditions there is a chain of ideals
$\mathrm{co}(R)\subseteq b(\omega)\subseteq\operatorname{\mathrm{tr}}(\omega)\subseteq\mathfrak{m}$
where $\mathrm{co}(R)$ is the conductor of $R$, provided that $R$ is not Gorenstein (in which case we have $b(\omega)=\operatorname{\mathrm{tr}}(\omega)=R$ ). The three extreme cases have been studied:
  1. (1)
    $R$ is a nearly Gorenstein ring if and only if the equality $\operatorname{\mathrm{tr}}(\omega)=\mathfrak{m}$ holds.
  2. (2)
    $R$ is an almost Gorenstein ring if and only if the equality $b(\omega)=\mathfrak{m}$ holds.
  3. (3)
    $R$ is a far-flung Gorenstein ring if and only if the equality $\operatorname{\mathrm{tr}}(\omega)=\mathrm{co}(R)$ holds.
In this paper, we are treating the missing case: the case where the equality $b(\omega)=\mathrm{co}(R)$ holds. We say that such a ring has minimal canonical conductor. By definition, the class of such rings contain the class of far-flung Gorenstein rings. Under some conditions, they coincide. For example, if $R$ has multiplicity 3, then $R$ has minimal canonical conductor if and only if it is far-flung Gorenstein. But in general the class of rings with minimal canonical conductor is much larger. In Section 3, we provide an example of a ring with minimal canonical conductor which is nearly Gorenstein.
In Section 2, we recall some preliminaries and prove equivalent conditions for having minimal conductor for an analytically unramified one dimensional local domain. Our main result in this section is the following.
Theorem A (Theorem 2.11).
Let $R$ be a one dimensional singular local domain which is analytically unramified. Let $\omega$ be a canonical ideal. Then, the following are equivalent.
  1. (1)
    The blowup algebra $B(\omega)$ is the normalisation $\overline{R}$.
  2. (2)
    There is an equality $b(\omega)=\mathrm{co}(R)$.
  3. (3)
    For any $\omega$-Ulrich module $M$, we have an isomorphism $M\cong\overline{R}^{\oplus r}$ for some $r$.
  4. (4)
    There exists an $n$ such that $\omega^{n}\cong\mathrm{co}(R)$.
In Section 3, we apply this theorem to numerical semigroup rings and classify numerical semigroup rings with minimal multiplicity.
Theorem B.
Let $H$ be a numerical semigroup with Frobenius number $f$. Then, the ring $k[[H]]$ has minimal canonical conductor if and only if f1f-1 is a quasi-Frobenius number. In particular, if $H$ is minimally generated by $n<a_{2}<\ldots<a_{n}$, then the ring $k[[H]]$ has minimal canonical conductor if and only if $a_{n}=a_{n-1}+1$.
Acknowledgements
I would like to thank Alfred Geroldinger, Alessio Moscariello and Shinya Kumashiro for valuable conversations and helpful comments.
2. Preliminaries and the main theorem
We start this section with some conventions and general definitions. Even though we give some definitions in a larger generality, we will always work with one dimensional Noetherian local domains throughout this paper and all modules will be finitely generated. Hence, our rings will be Cohen-Macaulay local rings and they will be generically Gorenstein. So, they will possess a canonical module.
We start with recalling some definitions.
Definition 2.1.
Let $R$ be a commutative Noetherian local ring .
  1. (1)
    Let $M$ be a finitely generated $R$-module. Then, the trace ideal $\operatorname{\mathrm{tr}}(M)$ of $M$ is the ideal of $R$ generated by the elements of the form $f(m)$ where $f\in M^{*}$ and $m\in M$.
  2. (2)
    Let $Q(R)$ be the total quotient ring of $R$. By $\overline{R}$, we denote the normalisation of $R$ inside $Q(R)$. That, is $\overline{R}$ is the integral closure of $R$ inside its total quotient rings. Then, the conductor ideal is defined as
    $\mathrm{co}(R)=\{r\in\overline{R}\colon r\overline{R}\subseteq R\}.$
    More generally, for any birational extension $R\subseteq S$, one can define
    $\mathrm{co}(R,S)=\{r\in S\colon rS\subseteq R\}.$
Remark 2.2.
Conductor ideals can be seen as trace ideals. This follows from [KT19, Proposition 2.4] and has been carefully treated in [HHS21, Appendix]. More precisely, one has $\mathrm{co}(R,S)=\operatorname{\mathrm{tr}}(S)$.
Remark 2.3.
Let $R$ be a one dimensional local domain which is analytically unramified. Assume that a finitely generated $R$-module has a nonzero trace ideal $\operatorname{\mathrm{tr}}(M)$ with a principal reduction. Then, there exists an inclusion $\mathrm{co}(R)\subseteq\operatorname{\mathrm{tr}}(M)$. When $R$ has infinite residue field, the principal reduction condition is satisfied automatically.
The following lemma was established by Faber [Fab20, Theorem 3.5] in the reduced case and more generally by Dao-Maitra-Sridhar [DMS23, Theorem 2.9].
Lemma 2.4.
Let $R$ be a commutative Noetherian ring, $S$ a finite birational extension and $M$ a finitely generated $R$-module. Consider the two statements:
  1. (1)
    $M$ is a module over $S$.
  2. (2)
    There is an inclusion $\operatorname{\mathrm{tr}}(M)\subseteq\mathrm{co}(R,S)$.
Then, (1) implies (2). When $M$ is reflexive, (2) implies (1), as well.
This gives us the following lemma about reflexive modules with minimal traces.
Lemma 2.5.
Let $R$ be a one dimensional local domain and assume that it is analytically unramified and $M$ a reflexive $R$-module of constant rank $r>0$. If there exists an equality $\operatorname{\mathrm{tr}}(M)=\mathrm{co}(R)$, then we have an isomorphism
$M\cong\overline{R}^{\oplus r}.$
Proof.
Let $M$ be a reflexive $R$-module with $\operatorname{\mathrm{tr}}(M)=\mathrm{co}(R)$. Then by Lemma 2.4, we know that $M$ is a module over $\overline{R}$. Since $M$ is reflexive, it is a maximal Cohen-Macaulay $R$-module and hence a maximal Cohen-Macaulay $\overline{R}$-module. Since $R$ is a one dimensional local domain, we know that $\overline{R}$ is a regular local ring. Hence, any maximal Cohen-Macaulay module over $\overline{R}$ is a direct sum of copies of $\overline{R}$. This finishes the proof. ∎
Definition 2.6.
Let $R$ be a commutative Noetherian ring and $I$ a regular ideal. Then, there is a filtration of endomorphism algebras
$R\subseteq I\colon I\subseteq I^{2}\colon I^{2}\subseteq\cdots I^{n}\colon I^{n}\subseteq\cdots\subseteq Q(A).$
We put
$R^{I}=\bigcup_{n\geq 0}I^{n}\colon I^{n}$
and denote by $b(I)$ the conductor ideal $\mathrm{co}(R,R^{I})$.
Remark 2.7.
Assume that $R$ is a commutative Noetherian ring of dimension one and $I$ a regular ideal.
  1. (1)
    The ring $R^{I}$ coincides with the blowup algebra $B(I)$ at $I$. This was proved by Lipman in his influencial work [Lip71] on Arf rings.
  2. (2)
    The ideal $b(I)$ equals the trace ideal $\operatorname{\mathrm{tr}}_{R}(R^{I})$.
Next, we will talk about $I$-Ulrich modules introduced by Dao-Maitra-Sridhar [DMS23]. The following definition is an equivalent condition for being an $I$-Ulrich module [DMS23, Theorem 4.6].
Definition 2.8.
Let $R$ be a one dimensional Cohen-Macaulay local ring and $I$ a regular ideal. Then a maximal Cohen-Macaulay $R$-module is called $I$-Ulrich if there is an isomorphism $IM\cong M$.
Remark 2.9.
Let $R$ be a one dimensional Cohen-Macaulay local ring and $I$ be a regular ideal.
  1. (1)
    A module $M$ is $I$-Ulrich if and only if $M\in\mathrm{CM}(B(I))$ [DMS23, Theorem 4.6]. This has the consequence that if $M$ is an $I$-Ulrich module, then one has an inclusion $\operatorname{\mathrm{tr}}(M)\subseteq b(I)$ and the converse holds when $M$ is reflexive.
  2. (2)
    If $M$ is an $I$-Ulrich module, then for any $X\in\mathrm{CM}(R)$, the module $\mathrm{Hom}_{R}(M,X)$ is also an $I$-Ulrich module [DMS23, Lemma 4.15].
  3. (3)
    If $R$ is analytically unramified, then the normalisation $\overline{R}$ is $I$-Ulrich.
Recall that a Cohen-Macaulay local ring $R$ is generically Gorenstein if for every minimal prime $p$ of $R$ the local ring $R_{p}$ is Gorenstein. In this case, the ring $R$ admits a canonical module $\omega$. In fact, the generically Gorenstein condition is equivalent to saying that the canonical module $\omega$ is isomorphic to an ideal. When we take $\omega$ to be an ideal we will refer to it as a canonical ideal.
Remark 2.10.
Let $R$ be a one dimensional generically Gorenstein local ring with canonical ideal $\omega$.
  1. (1)
    There is an equality $\mathrm{ref}(R)=\Omega\mathrm{CM}(R)$. That is, an $R$-module $M$ is reflexive if and only if it is the syzygy of a maximal Cohen-Macaulay module.
  2. (2)
    An $R$-module $M$ is $\omega$-Ulrich if and only if there is an isomorphism $M^{*}\cong DM$ [DMS23, Corollary 4.27].
  3. (3)
    For $n\gg 0$, the ideal $\omega^{n}$ is $\omega$-Ulrich.
We are now ready to state and prove the main theorem of this section.
Theorem 2.11.
Let $R$ be a one dimensional singular local domain which is analytically unramified. Let $\omega$ be a canonical ideal. Then, the following are equivalent.
  1. (1)
    The blowup algebra $B(\omega)$ is the normalisation $\overline{R}$.
  2. (2)
    There is an equality $b(\omega)=\mathrm{co}(R)$.
  3. (3)
    For any $\omega$-Ulrich module $M$, we have an isomorphism $M\cong\overline{R}^{\oplus r}$ for some $r$.
  4. (4)
    There exists an $n$ such that $\omega^{n}\cong\mathrm{co}(R)$.
Proof.
If two modules are isomorphic, then they have the same trace ideal which shows the implication $(1)\implies(2)$. If $M$ is an $\omega$-Ulrich module, then we have
$\mathrm{co}(R)\subseteq\operatorname{\mathrm{tr}}(M)\subseteq b(\omega)=\mathrm{co}(R).$
With this, Lemma 2.5 shows the implication $(2)\implies(3)$. Now assume $(3)$. By Remark 2.10(3), for $n\gg 0$, we know that $\omega^{n}$ is $\omega$-Ulrich. Hence, $\omega^{n}\cong\overline{R}^{\oplus r}\cong\mathrm{co}(R)^{\oplus r}$ for some $r$. As $\omega^{n}$ has rank 1, $r$ must be $1$ which proves $(4)$. Now, assume $(4)$. Then, we have equalities $B(\omega)=B(\omega^{n})=B(\mathrm{co}(R))=\overline{R}$. This implies $(1)$ showing that all conditions are equivalent. ∎
We end this section with our definition and a remark on relationship between far-flung Gorenstein rings.
Definition 2.12.
Let $R$ be a one dimensional singular local domain with canonical ideal $\omega$. Assume that $R$ is analytically unramified. We say that $R$ has minimal canonical conductor if the equivalent conditions of Theorem 2.11 hold for $R$.
Example 2.13.
Let $n\geq 3$ be a positive integer and $R=k[[t^{n},t^{n+1},\ldots,t^{2n-1}]]$. Then, one has equalities
$\mathrm{co}(R)=b(\omega)=\operatorname{\mathrm{tr}}(\omega)=\mathfrak{m}$
and $R$ has minimal canonical conductor and the canonical conductor is equal to the maximal ideal. Rings with this property have been considered by several authors in the last decade. For instance, they appear in Faber’s work [Fab20] as one-step normal rings. In the work of Dao-Dey-Dutta [DDD22], such rings are exactly one dimensional Ulrich split rings. In the language of Herzog-Kumashiro-Stamate [HKS23], they are far-flung Gorenstein and nearly Gorenstein. Recently, such rings also appeared in the work of Geroldinger-Yan-Zhang as conductor domains [GYQ].
Remark 2.14.
As we always have the inclusions $\mathrm{co}(R)\subseteq b(\omega)\subseteq\operatorname{\mathrm{tr}}(\omega)$ for a one dimensional analytically unramified local domain, the condition $\operatorname{\mathrm{tr}}(\omega)=\mathrm{co}(R)$ implies the condition $b(\omega)=\mathrm{co}(R)$. Hence, far-flung Gorenstein rings have minimal canonical conductor. However, not all rings with minimal canonical conductor are far-flung Gorenstein. We will see examples in the next section. On the other hand, we have the next proposition which shows when the two classes agree.
Before stating the proposition, we need a definition due to Kumashiro [Kum22].
Definition 2.15.
Let $R$ be a generically Gorenstein Cohen-Macaulay local ring. The infimum of all nonnegative integers $n$ such that there exists a canonical ideal $\omega$ and an almost reduction $(a)$ of $\omega$ with $\omega^{n+1}=a\omega^{n}$ is called the canonical reduction number of $R$ and is denoted by $\mathrm{can.red}(R)$.
Proposition 2.16.
For a one dimensional local domain, the following are equivalent.
  1. (1)
    $R$ has minimal canonical conductor and canonical reduction number 2.
  2. (2)
    $R$ is far-flung Gorenstein.
Proof.
The inclusion $b(\omega)\subseteq\operatorname{\mathrm{tr}}(\omega)$ is an equality if and only if there exists an isomorphism $\operatorname{\mathrm{tr}}(\omega)\cong\omega^{*}$ by [DL24, Proposition 4.7]. And such an isomorphism holds if and only if $R$ has canonical reduction number at most 2 by [Kum22, Theorem 2.13]. This shows that if $R$ has minimal canonical conductor and canonical reduction number 2, then it is far-flung Gorenstein. On the other hand, as mentioned above all far-flung Gorenstein rings have minimal canonical conductor. Moreover, by [HKS23, Theorem 2.5], we see that $R$ has canonical reduction number 2. This finishes the proof. ∎
We finish this section with an observation on the endomorphism ring $\mathfrak{m}\colon\mathfrak{m}$. This ring is often useful when one studies reflexive modules over $R$ and is far-flung Gorenstein when $R$. The latter assertion was proved by Herzog-Kumashiro-Stamate in [HKS23, Theorem 3.3] and the proof of the proposition below works similarly by making use of Claim 1 in the proof of op. cit.
Proposition 2.17.
Let $(R,\mathfrak{m})$ be a one dimensional local domain. If $R$ has minimal canonical conductor, then the ring $B=\mathfrak{m}\colon\mathfrak{m}$ also has minimal canonical conductor.
3. Numerical semigroups with minimal canonical conductor
In this section, we classify numerical semigroup rings with minimal canonical conductor. By a numerical semigroup $H$, we mean a subsemigroup of $\mathbb{N}$ with $\mathbb{N}\setminus H$ finite. This means that there exists a largest number $f$ which does not belong to $H$. This $f$ is called the Frobenius number of $H$. It has the property that $f+h\in H$ for any $h\in H$ with $h\neq 0$. This property can also hold for other numbers which do not belong to $H$. We put
$\mathrm{QF}(H):=\{x\in\mathbb{N}\colon x+h\in H\text{ for all }h\in H\text{ with }h\neq 0\}$
and call an element of this set a quasi-Frobenius number. The Frobenius number $f$ is the largest quasi-Frobenius number.
Given a field $k$, we denote by $k[[H]]$ the subring of the polynomial ring $k[[t]]$ generated by $t^{h}$ such that $h\in H$. There exists a nice interplay between $H$ and $k[[H]]$. For instance, every numerical semigroup has a minimal set of generators. The number of minimal generators of $H$ coincides with the embedding dimension of $k[[H]]$. The smallest element in $H$ coincides with the multiplicity of $k[[H]]$. Hence, $k[[H]]$ is of minimal multiplicity if and only if $H$ is minimally generated by elements of the form $n<a_{2}<\cdots<a_{n}$.
We can also see the type of $R=k[[H]]$ from the semigroup data. It is given by the size of $\mathrm{QF}(H)$. More precisely, if $f$ is the Frobenius number of $H$, then the $R$-module
$\omega=\sum_{x\in\mathrm{QF}(H)}Rt^{f-x}$
is a canonical module which is fractional ideal sitting between $R$ and $k[[t]]=\overline{R}$.
Theorem 3.1.
Let $H$ be a numerical semigroup with Frobenius number $f$. Then, the ring $R=k[[H]]$ has minimal canonical conductor if and only if f1f-1 is a quasi-Frobenius number for $H$.
Proof.
If f1f-1 is a quasi-Frobenius number for $H$, then there exists a canonical fractional ideal $\omega=R+Rt+J$ where $J=R^{t_{1}}+\ldots+R^{t_{s}}$ for some $t_{1},\ldots,t_{s}$. Note that $\omega^{n}\subseteq k[[t]]=\overline{R}$ for any positive integer $n$. On the other hand, we have
$(R+R^{t})^{n}=R+Rt+Rt^{2}+\cdots+Rt^{n}\subseteq\omega^{n}.$
For large enough $n$ (we can choose $n\leq f$), we have $(R+Rt)^{n}=k[[t]]=\overline{R}$. Hence, we obtain an isomorphism $\omega^{n}\cong\overline{R}$.
If f1f-1 is not a quasi-Frobenius number, we the canonical fractional ideal is of the form $\omega=R+R^{t_{1}}+\cdots+R^{t_{s}}$ for some $t_{1},\ldots,t_{s}\geq 2$. Therefore, we have $t\notin\omega^{n}$ for any $n\geq 1$. This finishes the proof. ∎
Corollary 3.2.
Let $H$ be a numerical semigroup minimally generated by $n<a_{2}<\cdots<a_{n}$. Then, the ring $k[[H]]$ has minimal canonical conductor if and only if $a_{n}=a_{n-1}+1$.
Proof.
Since the minimal number of generators equals the smallest generator, $k[[H]]$ is of minimal multiplicity. Therefore, quasi-Frobenius numbers of $H$ are given by $a_{2}-n,a_{3}-n,\ldots,a_{n}-n$. Hence, the Frobenius number is $a_{n}-n$ and $a_{n}-n-1$ is a Frobenius number if and only if $a_{n-1}=a_{n}-1$. ∎
Example 3.3.
Let $N$ be the semigroup generated by $n,n+1,n^{2}-n-1$ with $n\geq 3$ and $R=k[[N]]$ be the local semigroup ring generated by $N$. This ring was studied in [Kum22, Example 3.13]. It has a canonical fractional ideal $\omega=R+Rt$, it is nearly Gorenstein for all $n\geq 3$ and it is almost Gorenstein if and only if $n=3$. We see from Theorem 3.1 that
$\mathrm{co}(R)=b(\omega)\subseteq\operatorname{\mathrm{tr}}(\omega)=\mathfrak{m}.$
The inclusion $b(\omega)\subseteq\operatorname{\mathrm{tr}}(\omega)$ is strict if and only if $n>3$. Hence this ring, while being nearly Gorenstein, is far from being almost Gorenstein for $n>3$.